Minutes of nanoBPM meeting on 25 August, 2004

"Status of mover system", Y.Honda

(transparencies, 4 page, pdf, 402KB )

Three sets of movers have been fabricated. There was a repair of broken hinge in one of X-movers. A piece of hinge-part was fit in the X-mover, which was welded only on the surface. Also. warpage of the base plate will be repaired.

Vibrational resonance frequencies were measured to check their mechanical strength. The primary mode frequencies are listed below;
setXX'YY'xy
HzHzHzHz'HzHz
11502862864007004300
21203157204800
31003404500
, where x and y are active movers. Especially, the y-frequencies are consistent with calculations. All the frequencies seems to be high compared with ground motion. (Blank column will be filled soon.)

Dynamic range of movements was also measured by laser displacement sensors. The results are listed below;
setXX'YY'
μmμradμmμrad
1609360107
260underwayunderwayunderway
360underwayunderwayunderway
, where applied voltage was 150V for the X-mover.

Control system for the movers consists of ;

Most of them have been ordered to companies.

A rigid table (1 x 0.5m2) must be necessary near the Nano BPM system for the laser system surrounded by lead blocks. The location shall be checked at the ATF area.

"Status of KEK cavity BPM", Y. Inoue

(transparencies, xx page, pdf, xxxKB )

Electrical center of a c-band cavity BPM has been measured by antenna method. The preliminary results show that the center is shifted by 150um and 1mm in X and Y directions, respectively, from the mechanical center. Since the mechanical accurary is about 5 μm, the center should be coincident with this accuracy. One reason is a long antenna which could generate higher modes in feed-through. In order to avoid such problem, a metal bead (2mm diameter) was suspended into the cavity by a thin nylon string. Frequency changes have been observed at Δf/f=10-6 in the range of ±500μm. However, the 2mm &phi bead does not have enough accuracy. So, bead of larger diameter would be tested. A wire method may be another solution.

"Status of laser frame", Y. Higashi

Temperature system was fabricated, which has 20 channels with accuracy of 0.02 degree. The laser frame system (laser BPM, interferometer and temperature monitors) will be tested before the installation.

"Discussion on schedule"

We discussed schedule until the back-on of ATF beam on 18 October. The schedule is listed below;

However, we would like to bring forward the schedule by preference.

Next meeting will be held in 9 September, 2004, at 1:30pm-, 3-gokan 425.